Ron's age is two thirds of Mark's age. Mark's age is three fifths of Sara's age. Sara is 9 years older than Ron. How old is Mark ?
Answer by Maths68(1474)

Ron’s age=r
Mark’s age=m
Sara’s age=s
Sara is 9 years older than Ron means
r=s-9............(1)
Ron's age is two thirds of Mark's age
r=2/3(m)
r=2m/3........(2)
Mark's age is three fifths of Sara's age
m=3/5(s)
5m=3s
s=5m/3
put the value of r in equation (2)
s-9=2m/3..........(3)
now put the value of s in (3)
(5m/3)-9=2m/3
Multiply by 3 both sides of the equation
5m-27=2m
5m-2m=27
3m=27
m=27/3
m=9
Mark is 9 years old